OpenStudy (anonymous):

Find f ' (x) if f(x)= (x+2)/(x-1)

OpenStudy (anonymous):

use x/y differentiation rule.... d/dx(x/y)= (yx'-xy')/y^2

OpenStudy (anonymous):

\[\frac{(x+2)\prime \times(x-1)-(x-1)\prime(x+2)}{(x-1)^{2}}=\frac{1\times(x-1)-1(x-2)}{(x-1)^{2}}=\frac{1}{(x-1)^{2}}\]